OpenStudy (anonymous):

What is the value of w? 5(9 − 3) = w • 9 – 5 • 3

pooja195 (pooja195):

5(9−3)=w(9)−(5)(3) 30=9w+−15 30=9w−15 Flip the equation 9w−15+15=30+15 9w=45 Divide both sides by 9 45/9=?